# Historically Relevant Event Structuring for Temporal Knowledge Graph Reasoning

### v1

@@ 3.4.2. ConvGAT
3.5. Decoder
To achieve comprehensive predictions by leveraging entity and relation information from TKGs, $\mathsf{HisRes}$ adopts ConvTransE (Shang et al., 2019) as the score function, a widely used decoder in TKG reasoning, defined as follows:
(12) $p(o|s,r,t,\mathcal{G}_{0:t-1})=softmax(\mathbf{ConvTransE}(\textbf{s}^{\phi},\textbf{r}){\textbf{E}^{\phi}_{t}})\textnormal{,}$
(13) $\textbf{E}^{\phi}_{t}=\Theta^{\mathcal{H}}\textbf{E}^{\mathcal{H}}_{t}+(1-\Theta^{\mathcal{H}})\textbf{E}_{t}{,}$
(14) $\Theta^{\mathcal{H}}=\sigma(\textbf{W}_{8}\textbf{E}^{\mathcal{H}}_{t}+\textbf{b}_{gg})$
3.6. Learning Objective
Entity prediction in KGs can be regarded as a multi-classification task. Hence, $\mathsf{HisRes}$ employs Cross-Entropy as the loss function. Importantly, we optimize the relation prediction to enhance relation representations, akin to TiRGN (Li et al., 2022c) and LogCL (Chen et al., 2024), to further enhance the performance of entity prediction. In addition, we introduce a coefficient $\alpha$ to control the weight of the two tasks. The learning objective is formalized as follows:
(15) $\mathcal{L}$ $=\alpha\sum_{(s,r,t)\in\mathcal{Q}^{e}_{t}}y^{e}_{t}\log p(o|s,r,t,\mathcal{G}_{0:t-1})$
4. Experiments
@@
  • Q1: How does $\mathsf{HisRES}$ perform compared to state-of-the-art TKG reasoning studies?
@@
  • Q2: How do the multi-granularity evolutionary encoder and global relevance encoder affect the performance of $\mathsf{HisRES}$?
@@
  • Q3: How do the adaptive fusion of different entity representations and the updating of relations affect the performance of $\mathsf{HisRES}$?
  • Q4: How does the proposed superior ConvGAT affect the performance of $\mathsf{HisRES}$?
  • Q5: How does the performance fluctuation of $\mathsf{HisRES}$ with different settings?
@@
Table 3. TKG entity extrapolation results on ICEWS14s, ICEWS18, ICEWS05-15 and GDELT. The time-filtered MRR, H@1, H@3, and H@10 metrics are multiplied by 100. The best results are boldfaced and the second SOTAs are underlined.
@@
4.1. Experimental Setup
4.1.1. Datasets.
We conduct experiments on four real-world event-based TKGs extensively utilized in prior studies. Among them, three pertain to the Integrated Crisis Early Warning System (ICEWS) (Boschee et al., 2015): ICEWS14s, ICEWS18, ICEWS05-15. For comprehensive comparison and demonstrating the applicability and robustness of $\mathsf{HisRES}$, the Global Database of Events, Language, and Tone (GDELT) (Leetaru and Schrodt, 2013) dataset is also incorporated. Consistent with benchmarks (Li et al., 2021b; Chen et al., 2024), we split datasets into training (80%), validation (10%), and testing (10%) subsets based on their chronological sequences. Table 2 provides additional details on the statistical attributes of these datasets.
4.1.2. Baselines.
We compare our proposed $\mathsf{HisRES}$ with static KG reasoning methods (Yang et al., 2015; Trouillon et al., 2016; Dettmers et al., 2018; Shang et al., 2019; Sun et al., 2019) and the state-of-the-art TKG extrapolation models, which are shown below:
  • RE-NET (Jin et al., 2020) utilizes a combination of graph-based and sequential modeling to capture both temporal and structural interactions.
  • CyGNet (Zhu et al., 2021) constructs a historical vocabulary associated with all past events and rectifying predictions based on these events.
  • xERTE (Han et al., 2021) proposes a time-aware GAT that leverages temporal relations to explore subgraphs using temporal information.
  • RE-GCN (Li et al., 2021b) incorporates GNN and a recurrent unit to capture both the structural and temporal dependencies of TKGs, and introduces a static graph to augment initial embeddings.
  • CEN (Li et al., 2022a) employs a length-aware CNN within a curriculum learning and an online configuration to dynamically accommodate shifts in evolutionary patterns.
  • TiRGN (Li et al., 2022c) proposes time-aware decoders and incorporates a global historical vocabulary to predict based on whole history.
  • CENET (Xu et al., 2023) introduces contrastive learning into TKGs and designs a classifier to distinguish historical or global events.
  • RETIA (Liu et al., 2023) combines line graphs and captures the spatial-temporal patterns of entities and relations simultaneously.
  • RPC (Liang et al., 2023) incorporates a rule-based line graph for relations and mines graph structural correlations and periodic temporal interactions via two correspondence units.
  • LogCL (Chen et al., 2024) proposes an entity-aware attention encoder to capture crucial facts from local and global history.
4.1.3. Implementation Details.
The initial embedding dimension $d$ is configured to 200 for all datasets. The dropout rate of 0.2 is applied uniformly across all modules. The layers of GNNs are set to 2. Additionally, we utilize the Adam optimizer with a learning rate parameterized to 0.001. For the optimal historical length $l$ in multi-granularity evolutionary encoder, we conduct grid research according to the validation results, and the best value of $l$ is set to 9, 9, 10 and 7 for ICEWS14s, ICEWS18, ICEWS05-15 and GDELT, respectively. Noted that, to fair comparison, the task coefficient $\alpha$ is set to 0.7, and the static graph learning module is employed on ICEWS datasets similar to our baselines (Li et al., 2021b, 2022c; Liang et al., 2023; Chen et al., 2024). All baseline results are reported by previous works. We adopt a two-phase forward propagation strategy for the raw and inverse query sets similar to LogCL (Chen et al., 2024). All experiments are conducted on NVIDIA A800 GPUs.
4.1.4. Evaluation Metrics.
We report the widely-used time-aware filtered metrics (Han et al., 2021; Li et al., 2022c; Liang et al., 2023; Chen et al., 2024) (MRR and Hits@1,3,10) with percentage form in this paper.
4.2. Results (RQ1)
@@
The performance of $\mathsf{HisRES}$ and advanced baselines under time-aware filtered metrics is reported in Table 3.
@@
To comprehensively assess the performance of $\mathsf{HisRES}$, we initially compare it with static KG reasoning methods (i.e., DistMult, ComplEx, ConvE, ConvTransE, RotatE). As existing TKG baselines consistently outperform these methods across all benchmarks, underscoring the importance of considering temporal information and modeling evolutionary patterns within TKGs. $\mathsf{HisRES}$ achieves more progressive improvements with a 58.24% increase in MRR compared to ConvE. Furthermore, to provide valuable insights and compelling evidence addressing Q1, we compare $\mathsf{HisRES}$ with the best-performing TKG models. It demonstrates that $\mathsf{HisRES}$ is the state-of-the-art method, exhibiting a significant average improvement of 6.11%, 8.41%, 6.54% and 3.69% in time-filtered MRR, Hits@1, Hits@3 and Hits@10, respectively, across four event-based datasets.
@@
Specifically, CyGNet and CENET focus on statistical analysis of historically repetitive facts but tend to overlook capturing structural interactions among snapshots. In contrast, $\mathsf{HisRES}$ not only models the structural and temporal dependencies of recent snapshots, but also incorporates a more comprehensive understanding of the entire history with the proposed global relevance encoder. It is more effective than simply rectifying predicted results.
@@
In comparison to a series of approaches (Jin et al., 2020; Li et al., 2021b, 2022a, 2022c; Liu et al., 2023; Liang et al., 2023) that mainly focus on modeling each single snapshots from the most recent history, $\mathsf{HisRES}$ additionally considers the multi-granularity recent history. Concretely, we learn the interaction of each adjacent snapshot simultaneously (across time) while utilizing multi-hop links of GNNs to correlate the evolution of events intra-snapshot. Besides, we introduce a novel perspective that exclusively concentrates on relevant events throughout the entire history and further filters important links from them through ConvGAT. Although TiRGN considers global historical information, it still faces challenges similar to CyGNet, as it only employs a simple vector to represent global repetitive facts.
@@
Furthermore, $\mathsf{HisRES}$ surpasses LogCL which achieves the second-best results and also provides a global perspective on top of recent snapshot modeling. This outstanding performance stems from multi-granularity and considering a self-gating mechanism to adaptively merge entity representations from different historical periods. Meanwhile, a more efficient approach to handling globally relevant facts is also helpful. In detail, we focus solely on facts directly relevant to the queries, thus reducing the number of historical facts. Additionally, integrating attention calculations into ConvGAT assists in selecting high-impact links or facts.
@@
It should be noted that previous methods have not achieved significant improvements on GDELT due to the finer granularity of the time interval per snapshot (15 minutes), which makes GDELT more time-sensitive. $\mathsf{HisRES}$ provides an effective way of capturing multi-granularity features from both local evolution and global history while achieving a significant improvement of 11.92%, 15.44%, 13.55% and 9.40% in MRR, Hits@1, Hits@3, and Hits@10 on GDELT. This highlights the ability of $\mathsf{HisRES}$ to capture correlations within highly time-sensitive datasets, which have the typical characteristics of TKGs.
@@
Table 4. Ablation studies with time-filtered metrics on ICEWS14s and ICEWS18.
Model ICEWS14s ICEWS18
@@
$\mathsf{HisRES}$ 50.48 39.57 56.65 71.09 37.69 26.46 42.75 59.70
$\mathsf{HisRES}$-w/o-$\mathcal{G}$ 45.48 34.76 50.94 65.72 29.16 18.45 33.17 50.61
$\mathsf{HisRES}$-w/o-$\mathcal{G}^{\mathcal{H}}$ 41.83 31.49 47.01 61.74 31.55 21.53 35.41 51.48
$\mathsf{HisRES}$-w/o-MG 49.67 38.95 55.55 70.11 36.31 25.11 41.09 58.49
$\mathsf{HisRES}$-w/o-$\textit{SG}^{1}$ 50.04 39.34 55.86 70.28 37.08 25.76 42.07 59.39
$\mathsf{HisRES}$-w/o-$\textit{SG}^{2}$ 50.10 39.42 56.24 70.07 36.99 25.70 41.95 59.39
$\mathsf{HisRES}$-w/o-RU 50.17 39.37 56.17 70.38 36.99 25.79 41.79 59.12
$\mathsf{HisRES}$-w/-CompGCN 48.75 37.71 54.70 69.73 36.37 25.34 41.06 58.21
$\mathsf{HisRES}$-w/-RGAT 47.99 36.95 53.94 69.18 35.68 24.58 40.30 57.72
4.3. Ablation Studies (RQ2, RQ3 and RQ4)
To investigate the impact of different components within $\mathsf{HisRES}$ (Q2 and Q3), as well as the superiority of proposed ConvGAT (Q4), we conduct a series of ablation studies with ICEWS14s and ICEWS18 as representative datasets. The findings and results are summarized in Table 4.
  • The impact of different encoders (RQ2). To evaluate the indispensable local and global relevance structuring in TKG reasoning, we introduce two variants: $\mathsf{HisRES}$-w/o-$\mathcal{G}$ and $\mathsf{HisRES}$-w/o-$\mathcal{G}^{\mathcal{H}}$, indicating the removal of the multi-granularity evolutionary encoder and the global relevance encoder, respectively. The former highlights the importance of capturing evolutionary patterns from different granularities within the most recent snapshots. Conversely, the latter emphasizes the importance of additionally capturing global dependencies apart from local patterns, which notably enhances TKG reasoning.
@@ 4.3. Ablation Studies (RQ2, RQ3 and RQ4)
  • The impact of self-gating and relation updating (RQ3). To investigate the effectiveness of adaptively merging different entity representations of TKGs, we replace self-gating with a simple summation. The variants are denoted as $\mathsf{HisRES}$-w/o-$\textit{SG}^{1}$ and $\mathsf{HisRES}$-w/o-$\textit{SG}^{2}$, corresponding to the two-stage fusion (equation (8) and (13)) in our methodology. The decreased performance shown in the second part of Table 4 indicates that the self-gating mechanism is effective for either type of entity representation fusion. Additionally, our strategy of relation updating (represented as $\mathsf{HisRES}$-w/o-RU) achieves an average improvement of 1.26% in MRR, underscoring the importance of considering the evolution of relations during aggregations.
@@
  • To assess the superiority of the proposed ConvGAT, we replace it with two widely used GNN aggregators, CompGCN (Vashishth et al., 2020) and RGAT (Nathani et al., 2019). The resulting variants are denoted as $\mathsf{HisRES}$-w/-CompGCN and $\mathsf{HisRES}$-w/-RGAT, respectively, as shown in the third part of Table 4. Compared to the second-best result, the original $\mathsf{HisRES}$ achieves an average improvement of 3.59%, 4.68%, 3.84% and 2.26% in MRR, Hits@1, Hits@3 and Hits@10, respectively. It illustrates the significance of assigning different degrees of attention to various historical facts and underscores the efficiency of ConvGAT in calculating attention scores.
4.4. Sensitivity Analysis (RQ5)

### v2

@@ IV Experiments
  • Q1: How does HisRES perform compared to state-of-the-art TKG reasoning studies?
@@
  • Q2: How do the multi-granularity evolutionary encoder and global relevance encoder affect the performance of HisRES?
@@
  • Q3: How do the adaptive fusion of different entity representations and the updating of relations affect the performance of HisRES?
  • Q4: How does the proposed superior ConvGAT affect the performance of HisRES?
  • Q5: How does the performance fluctuation of HisRES with different settings?
IV-A Experimental Settings
IV-A1 Datasets
We evaluate our approach on four real-world event-based TKGs extensively utilized in prior studies. Three datasets are derived from the Integrated Crisis Early Warning System (ICEWS) [17] with daily temporal granularity: ICEWS14s, ICEWS18, and ICEWS05-15. ICEWS14s and ICEWS18 both cover one-year periods, but ICEWS18 contains approximately three times more entities and five times more facts per snapshot than ICEWS14s. ICEWS05-15 encompasses an extended timespan of 11 years. To demonstrate the broad applicability, scalability, and robustness of HisRES through comprehensive evaluation, we also include the Global Database of Events, Language, and Tone (GDELT) [43]. This dataset features ultra-fine-grained temporal resolution (15-minute snapshots) and high event density (2,278,405 facts within a one-month period). Performance gains on this challenging large-scale dataset provide a more rigorous validation of the model’s effectiveness. Consistent with benchmarks [6, 16], all methods adhere to the split strategy that divides datasets into training (80%), validation (10%), and testing (10%) subsets based on their chronological sequences. Table III provides additional details on the statistical attributes of these datasets.
Notably, results on some benchmarks such as WIKI and YAGO are not reported. These datasets have large time granularity (one year per snapshot) and long time spans (232 and 189 years, respectively). Such coarse-grained, long-term datasets may be less time-sensitive and less suitable for TKG applications, as knowledge spanning over a century may not be as relevant. Consequently, most recent SOTAs (e.g., L2TKG and LogCL) do not report results on these datasets.
IV-A2 Baselines
We compare our proposed HisRES with static KG reasoning methods [22, 23, 20, 25, 24, 26, 27, 29] and the state-of-the-art TKG extrapolation models, including RE-NET [10], CyGNet [8], xERTE [34], RE-GCN [6], EvoKG [44], CEN [11], TiRGN [7], CENET [9], RETIA [12], RPC [13], and LogCL [16]. HGLS and L2TKG are not reported due to different experimental settings. The details of these extrapolation methods are shown below:
  • RE-NET [10] utilizes a combination of graph and sequential modeling to capture temporal and structural interactions.
  • CyGNet [8] constructs a historical vocabulary associated with all past events to rectify predictions.
  • xERTE [34] proposes a time-aware GAT that leverages temporal relations to explore subgraphs using temporal information.
  • RE-GCN [6] incorporates GNN and a recurrent unit to capture both the structural and temporal dependencies of TKGs, and introduces a static graph to augment initial embeddings.
  • CEN [11] employs a length-aware CNN within a curriculum learning and an online configuration to dynamically accommodate shifts in evolutionary patterns.
  • TiRGN [7] proposes time-aware decoders and incorporates a global historical vocabulary to predict based on whole history.
  • CENET [9] introduces contrastive learning into TKGs and designs a classifier to distinguish historical or global events.
  • RETIA [12] combines line graphs and captures the spatial-temporal patterns of entities and relations simultaneously.
  • RPC [13] incorporates a rule-based line graph for relations and mines graph structural correlations and periodic temporal interactions via two correspondence units.
  • LogCL [16] proposes an entity-aware attention encoder to capture crucial facts from local and global history.
IV-A3 Implementation Details
The initial embedding dimension $d$ and training epoch are configured to 200 and 30 for all datasets, respectively. The dropout rate of 0.2 is applied uniformly across all modules. The granularity span $\omega$ and hidden layers of GNNs are set to 2 for all datasets. Additionally, we utilize the Adam optimizer with a learning rate parameterized to 0.001. For the optimal historical length $l$ in multi-granularity evolutionary encoder, we conduct grid research according to the validation results, and the best value of $l$ is set to 9, 9, 10 and 7 for ICEWS14s, ICEWS18, ICEWS05-15 and GDELT, respectively. For fair comparison, the task coefficient $\alpha$ is empirically set to 0.7, and the static graph is employed on ICEWS datasets similar to our baselines [6, 7, 13, 16]. All baseline results are reported by previous works. For the global relevance encoder, we adopt a two-phase forward propagation strategy for both raw and inverse query sets, similar to LogCL [16]. This approach enables our model to construct globally relevant graphs for the raw and inverse query sets separately, thereby preventing potential data leakage. All experiments are conducted on NVIDIA A800 GPUs.
IV-A4 Evaluation Metrics
We report the widely-used time-aware filtered metrics: Mean Reciprocal Rank (MRR) and Hits@{1,3,10} [34, 7, 13, 16], with results presented in percentage form.
@@ IV-A4 Evaluation Metrics
TABLE IV: TKG entity extrapolation results on ICEWS14s, ICEWS18, ICEWS05-15 and GDELT. The time-filtered MRR, H@1, H@3, and H@10 metrics are multiplied by 100. The best results are boldfaced and the second SOTAs are underlined.
@@
IV-B Main Results (RQ1)
@@
The performance of HisRES and advanced baselines under time-aware filtered metrics is reported in Table IV.
@@
To comprehensively assess the performance of HisRES, we initially compare it with static KG reasoning methods (i.e., DistMult, ComplEx, ConvE, ConvTransE, RotatE, GCN, R-GCN, GAT). As existing TKG baselines consistently outperform these methods across all benchmarks, underscoring the importance of considering temporal information and modeling evolutionary patterns within TKGs. Quantitatively, HisRES achieves more progressive improvements in MRR of 64.30% over ComplEx (translation-based), 58.24% over ConvE (convolution-based), and 57.41% over GCN (GNN-based). Furthermore, to provide valuable insights and compelling evidence addressing Q1, we compare HisRES with the best-performing TKG models. It demonstrates that HisRES is the state-of-the-art method, exhibiting a significant average improvement of 6.11%, 8.41%, 6.54% and 3.69% in time-filtered MRR, Hits@1, Hits@3 and Hits@10, respectively, across four event-based datasets.
@@
Specifically, CyGNet and CENET focus on statistical analysis of historically repetitive facts but tend to overlook capturing structural interactions among snapshots. In contrast, HisRES not only models the structural and temporal dependencies of recent snapshots, but also incorporates a more comprehensive understanding of the entire history with the proposed global relevance encoder. It is more effective than simply rectifying predicted results.
@@
In comparison to a series of approaches [10, 6, 11, 7, 12, 13] that mainly focus on modeling each single snapshots from the most recent history, HisRES additionally considers the multi-granularity recent history. Concretely, we learn the interaction of each adjacent snapshot simultaneously (across time) while utilizing multi-hop links of GNNs to correlate the evolution of events intra-snapshot. Besides, we introduce a novel perspective that exclusively concentrates on relevant events throughout the entire history and further filters important links from them through ConvGAT. Although TiRGN considers global historical information, it still faces challenges similar to CyGNet, as it only employs a simple vector to represent global repetitive facts.
@@
Furthermore, HisRES surpasses LogCL which achieves the second-best results and also provides a global perspective on top of recent snapshot modeling. This outstanding performance stems from multi-granularity and considering a self-gating mechanism to adaptively merge entity representations from different historical periods. Meanwhile, a more efficient approach to handling globally relevant facts is also helpful. In detail, we focus solely on facts directly relevant to the queries, thus reducing the number of historical facts. Additionally, integrating attention calculations into ConvGAT assists in selecting high-impact links or facts.
@@
Notably, previous methods have failed to achieve significant improvements on GDELT due to the finer temporal granularity per snapshot (15 minutes), which makes GDELT more time-sensitive. HisRES provides an effective way of capturing multi-granularity features from both local evolution and global history while achieving a significant improvement of 11.92%, 15.44%, 13.55% and 9.40% in MRR, Hits@1, Hits@3, and Hits@10 on GDELT. This highlights the ability of HisRES to capture correlations within highly time-sensitive datasets, which have the typical characteristics of TKGs.
IV-C Ablation Studies (RQ2, RQ3 and RQ4)
To investigate the impact of different encoders (Q2) and components (Q3) within HisRES, as well as the superiority of proposed ConvGAT (Q4), we conduct a series of ablation studies with ICEWS14s and ICEWS18 as representative datasets. The findings and results are summarized in Table V.
IV-C1 The Impact of Different Encoders (RQ2)
To evaluate the indispensable local and global relevance structuring in TKG reasoning, we introduce two variants: HisRES-w/o-$\mathcal{G}$ and HisRES-w/o-$\mathcal{G}^{\mathcal{H}}$, indicating the removal of the multi-granularity evolutionary encoder and the global relevance encoder, respectively. The former highlights the importance of capturing evolutionary patterns from different granularities within the most recent snapshots. Conversely, the latter emphasizes the importance of additionally capturing global dependencies apart from local patterns, which notably enhances TKG reasoning. Additionally, HisRES-w/o-MG indicates that our multi-granularity evolutionary encoder only uses intra-snapshot modeling. The resulting deterioration in scores demonstrates the importance of considering both intra- and inter-snapshot information simultaneously.
@@ IV-C1 The Impact of Different Encoders (RQ2)
TABLE V: Ablation studies with time-filtered metrics on ICEWS14s and ICEWS18.
Variant Model ICEWS14s ICEWS18
@@
HisRES 50.48 39.57 56.65 71.09 37.69 26.46 42.75 59.70
HisRES-w/o-$\mathcal{G}$ 45.48 34.76 50.94 65.72 29.16 18.45 33.17 50.61
HisRES-w/o-$\mathcal{G}^{\mathcal{H}}$ 41.83 31.49 47.01 61.74 31.55 21.53 35.41 51.48
HisRES-w/o-MG 49.67 38.95 55.55 70.11 36.31 25.11 41.09 58.49
HisRES-w/o-$\textit{SG}^{1}$ 50.04 39.34 55.86 70.28 37.08 25.76 42.07 59.39
HisRES-w/o-$\textit{SG}^{2}$ 50.10 39.42 56.24 70.07 36.99 25.70 41.95 59.39
HisRES-w/o-TE 50.08 39.47 55.83 70.54 37.25 26.10 42.14 59.24
HisRES-w/o-RU 50.17 39.37 56.17 70.38 36.99 25.79 41.79 59.12
HisRES-w/-CompGCN 48.75 37.71 54.70 69.73 36.37 25.34 41.06 58.21
HisRES-w/-RGAT 47.99 36.95 53.94 69.18 35.68 24.58 40.30 57.72
IV-C2 The Impact of Different Components (RQ3)
@@
To investigate the effectiveness of adaptively merging different entity representations, we replace self-gating at two parts (corresponding to equations (8) and (13)) with simple summation. These variants are denoted as HisRES-w/o-$\textit{SG}^{1}$ and HisRES-w/o-$\textit{SG}^{2}$. The decreased performance shown in the second part of Table V indicates that the self-gating mechanism is effective for either type of entity representation fusion. Subsequently, HisRES-w/o-TE (corresponding to equations (1) and (2)) exhibits a 0.98% decrease in MRR, highlighting the impact of considering the temporal gap in our multi-granularity evolutionary encoder. Additionally, compared to HisRES-w/o-RU (corresponding to equations (4) and (5)), our strategy of relation updating achieves an average improvement of 1.26% in MRR, underscoring the importance of considering the evolution of relations during aggregations.
IV-C3 The Impact of Different GNNs (RQ4)
@@
To assess the superiority of the proposed ConvGAT, we replace it with two widely used GNN aggregators, CompGCN [28] and RGAT [30]. The resulting variants are denoted as HisRES-w/-CompGCN and HisRES-w/-RGAT, respectively, as shown in the third part of Table V. Compared to the second-best result, HisRES achieves an average improvement of 3.59%, 4.68%, 3.84% and 2.26% in MRR, Hits@1, Hits@3 and Hits@10, respectively. It illustrates the significance of assigning different degrees of attention to various historical facts and underscores the efficiency of ConvGAT in calculating attention scores. Moreover, to demonstrate ConvGAT’s ability to focus on crucial historical cues, we conduct an additional analysis of attention weight distributions, as detailed in Section IV-H.
IV-D Sensitivity Analysis (RQ5)